OpenStudy (anonymous):

if gcd (a 42)= 1 show that 168=3*7*8 divide a^6-1

OpenStudy (anonymous):

since gcd (a, 42) = 1, gcd(a, 2) = 1 gcd(a, 3) = 1 gcd(a, 7) = 1

OpenStudy (anonymous):

hmm...now solve for (a,2)=1 only

OpenStudy (anonymous):

sicne 2, 3, 7 are relatively prime first prove 42 | a^6-1 using fermat's theorem

OpenStudy (anonymous):

and then??? how can we convert mod 42 to mod 168???

OpenStudy (anonymous):

good question , let me think a bit, i did similar problem few days back. let me see if i can pull it up

OpenStudy (anonymous):

ok i m waiting

OpenStudy (anonymous):

look at the LAST REPLY of ikram

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz explain this step 2|a^4-1 for a odd (gcd(a,2)=1) (2k+1)^4-1 = (4(...)+1)-1=4(....) thus 4| a^4-1...............eq2

OpenStudy (anonymous):

this is last reply of ikram

OpenStudy (anonymous):

since gcd(a, 2) = 1 a = 1 (mod 2)

OpenStudy (anonymous):

so, a^4 = 1 (mod 2) 2 | a^4-1 fine so far ?

OpenStudy (anonymous):

yes...next?

OpenStudy (anonymous):

then we see that, since gcd(a, 2) = 1, a must be odd.

OpenStudy (anonymous):

a = 2k + 1

OpenStudy (anonymous):

next, consider : a^4-1 (2k+1)^4 - 1 you can write it as 4(some stuff) so 4 | a^4-1

OpenStudy (anonymous):

that means, gcd(a, 4) = 1

OpenStudy (anonymous):

key point is : (2k+1)^4 - 1 is divisible by 4 always

OpenStudy (anonymous):

we can say...if any odd no is relatively prime to 2...then it is also relatively prime to every multiple of 2...right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

we can say...if any odd no is relatively prime to 2...then it is also relatively prime to every POWER of 2...

OpenStudy (anonymous):

i think u mean that right ? then it looks right

OpenStudy (anonymous):

ok like 9 and 12 are not relatively prime

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

ok...thanks a lot.... now i try to solve... if i get any confusion i need ur help again...

OpenStudy (anonymous):

okay, good luck :)

OpenStudy (anonymous):

@ beccaboo022a in my que a problem arise (a,2)=1 a=1 mod 2 a^6=1 mod 2 2Ia^6-1 as a is odd so let a=2k+1 but a^6-1=(2k+1)^6-1 is not divisible by 8....:(

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

to show 8 | a^6-1 : since 2 | a^6-1 thus, a is odd say a = 2k + 1 consider : a^2 (2k+1)^2 = 4k^2 + 4k + 1 = 4k(k+1) + 1 = 8m + 1

OpenStudy (anonymous):

thus, a^2 = 1 mod 8

OpenStudy (anonymous):

see if that makes some sense. we used this :- k(k+1) = 2m

OpenStudy (anonymous):

ok i understand a new method:(

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its old method oly :) just a different way i think

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

so every odd number square is of form 8m + 1